FEL Upgrade Project Weekly Brief
May 28 - June 1, 2001

We began the process of recertifying the Laser Safety System (LSS) for the start-up of the summer operations

The project monthly report for April was distributed to the ONR and DOE program offices this week.

We made final preparations for our participation in next week’s (June 5-6) JTO-sponsored workshop on
“Free-Electron Lasers  for Naval Applications” which is being held at the Ramada Hotel in Newport News.
Over 120 participants are expected from the Navy, industry and the government labs.

On Thursday May 31, three visitors from DARPA (Tim Grayson, Gail Heim and Sheldon Meth)
visited Jefferson Lab for  a tour and briefing on the FEL program.

WBS 3 (Beam Physics):
A review of the QX quadrupole systematic multipole specification was conducted.  As the bore is larger than that
in the Demo, the specification can be relaxed from 1 ppt to 3 ppt without untoward performance effects.  The
prototype magnet meets the revised specification. The review is documentented in JLAB-TN-01-027, available
on line at http://www.jlab.org/~douglas/FELupgrade/technote/JLABTN01027.pdf

WBS 4 (Injector):
Received the 6" RF shielded gate valve from VAT. Submitted the 10 to 6 MV/m spacer to the shop for
fabrication.  Signed off the remaining shield door test drawings.  Reviewing the stalk drawings for sign off.

WBS 5 (SRF):
The weld development for the HOM couplers are complete with the first two prototypes completed.

Cavity tests have shown a sloping Qo value on cavity JL005 with full chemistry.  This problem has lead to several
tests to determine the cause of the slope.  So far tests have indicated that the problem is not the indium seal and
the next test is to heat-treat cavity JL005 to 800C for 6 hours to determine if niobium hydride precipitation (Qo
disease)is the problem.

Flange leaks have developed on the FPC flange and testing to resolve the issue has started.

WBS 8 (Instrumentation):
Getting ready for the user run has been a priority this week.  All of the LSS Lab Interface Box power supplies
were upgraded to handle increased current caused by more equipment in the labs.  The low level certifications
went well, there was only one minor hitch with the series/parallel hutch bypass contacts on the optical transport

OBPM Picomotor controller chassis in system check-out and diagnostic software made. Many of the zero &
tracking functions of the OBPM system have been tested using the drive laser. We still only have 2 of the
Molectron sensors in house; the remaining prototype boards are complete with the exception of those sensors
which are due in next week.

Requests for bids on assembly of the new Ion Pump Power Supplies went out late last week.  One bid has been
received so far.

The new silicon viewers have been tested with beam and work very well, a decision whether to attenuate the
signal will be made Friday.

Work continues on the user power supplies and the web site upgrade.

Drawings were completed this week for the Front PCB in the Ion Pump Power Supply chassis; the schematic of
the PCB in the Beamviewer Solenoid Control Box; and the Front Panel Fabrication of the VME Timing Circuit.

Drawings in EECAD:  IPPS Main Board Schematic; MPS System Drawing; Analog Differential Driver Board
Schematic; Pyro-Electric OBPM Buffer/Amplifier; Picomotor Relay Chassis Wiring Diagram; VME Timing
Module Schematic, Fabrication and Assembly. VME programmable timing module front panel drawing
signed-off, PCB design/schematic proceeding. VME based MPS input card front panel layout and PCB design
progress.  VME Interface daughterboard interconnections are specified and the PCB design is progressing as

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o DULY Research, using their magnetic model, probed the sensitivity of their sliding shorts to the field clamps
    solution to uniform field integral of the small injector dipole. They found that changes of a mm keep the integral
    within specification.
o They sent the complete package of drawings of the small injector dipole (DV) to us for checking.
Arc Dipoles (GY, GX, GQ)
o AES was successful with new runs with the model of the 180 degree Dipole with reduced strength path length
    correctors.  The good field region width was expanded and the longitudinal field uniformity also improved
    significantly.  With a small change in the bulk field obtained by adjusting the Purcell Gap, the parallel offset would
    also fall into the sub mm values that are within David Douglas’ specifications.  Further work was halted until the
    current and bulk field were appropriately matched between  this magnet and the Arc Bend  (GX).
o They  are continuing design and magnetic modeling the Arc Bend Dipole (GX).  Their results showed a
    parabolic change in  the bulk field of the magnet, lowering by 7 parts in 10000 over the 0.6 meter central zone.
    Consultations with David Douglas and further efforts are on going to find out if this is acceptable, to see if we
    experienced similar results with the IR Demo and on determining just what Bulk Field really means when applied
    to a real magnet.
o Design efforts concentrated on introducing the lead configuration that allows powering the two sets of coil pairs
QX (3 inch quad)
o The coil manufacturing bids were received and the evaluation team given the papers to start their process.
o  The core procurement package remains out for bid.
o  One of the added end chamfer shim configurations to the prototype quadrupole showed promise in
    understanding the reduction of the n = 10 pole and another run with double the shim confirmed showed double
    the results.  We will now try a series of cuts using this understanding to reduce both the n = 6 pole and the n = 10
    pole.  David Douglas has found that he can live with the higher 3 part per thousand error of the original
    configuration but is encouraging us to get a lower value.
QT (Trim Quad)
o  Design work continued.
o  Robin Wines restudied the 2D design to understand further the generation of a  vertical corrector dipole set of
    windings.  She is now continuing to change the model to 3D design in TOSCA.
o  We anticipate a response today from the company that gave us the Dipole Steel bid so that we can place that
o  We are now assembling the attachments for the task definition to be used by an engineering service contractor
    for design of the arc chambers.
o  Jeff Karn gave us notice that he will be leaving the Lab on the 20th of June.  Robin Wines will take over
    direction of the magnet effort and Tom Hiatt will take over direction of Magnet Measurements.

WBS 11 (Optics):
The optical cavity assembly design has progressed in the following areas: The movement of the gimbal mounts
has been converted from linear translation stages to precision rails and ball bushings.  We decided to reduce
some complication in the framing by attaching the rails to the side of the vacuum enclosure.  The gimbal assembly
support structure, which also supports the linear actuators and lvdts, has been designed.  We received quotes
and engineering details on another UHV motorized actuator.  The actuator surpasses our specifications, and
should integrate easily into our design.  We are checking with at least two other vendors to confirm the pricing is
competitive.  We received a motorized actuator and will start outgassing studies next week.

The following work was done this week to support Operations:  Removed optics diverting the drive laser beam
to the OBPM test stand.  Aligned turning mirrors for the bypass transport to the carbon anotubes experiment.
We received refigured, broadband (5.0 - 6.6 um) outcouplers, and are inspecting them.

Preparations continued for the summer user run including check-out of diagnostics and LSS certification.  We
also continued working with various groups in the Lab to help with the re-certification of the ODH sensors.  At
this writing the Safety Systems group has identified a recalibration procedure that will allow the existing ODH
sensors to be recalibrated for the 18% oxygen concentration set-point.  This procedure will be attempted with
the ODH sensors in the FEL vault on Monday.

Privacy and Security Notice

Privacy and Security Notice